OpenStudy (anonymous):

Help Please n_n - what is the value of p? x=1/4 y^2

OpenStudy (anonymous):

its a parabola \[x=\frac{ 1 }{ 4 } y^2\]

OpenStudy (campbell_st):

well the standard form is \[4p(x - h) = (y - k)^2\] it appears the vertex is at the origin so you have \[4px = y^2\] so I'd multiply both sides of your equation by 4 so you can compare it to the standard form and then find p

OpenStudy (anonymous):

ok so i got 4x=y^2 anddd if i compare this to the above one p=1 ??? @campbell_st

OpenStudy (campbell_st):

that's what my answer would be

OpenStudy (anonymous):

ohhhhh okkkk thank you